OpenStudy (dannyboi):

Find the 20th term of the following sequence. -6, -4, -2, 0, ... 1)32 2)44 3)46

OpenStudy (dallascowboys88):

+2,+4,+6 just like on number line.negative on one side positive the other,but using the same numbers just switching signs.it's simple too do if you think about it.

OpenStudy (dallascowboys88):

A number line.

OpenStudy (dallascowboys88):

#1=Correct

OpenStudy (jango_in_dtown):

its an AP , first term=-6 and common difference =2 so tn=a+(n-1)d where tn is the nth term, a=-6, d=2. thus t20=-6+(20-1)*2

OpenStudy (jango_in_dtown):

=-6+38

OpenStudy (dannyboi):

thanx

OpenStudy (dallascowboys88):

exactly what i just said,just in different words.

OpenStudy (dallascowboys88):

np